OpenStudy (anonymous):

Sqrt(3)*cot[theta] -1=0 Need to solve for theta

OpenStudy (thomas5267):

\[\cot(\theta)=\frac{1}{\tan(\theta)}\]

OpenStudy (anonymous):

\[\cot \theta=\frac{ 1 }{ \sqrt{3} }\] \[\tan \theta=\frac{ \sqrt{3} }{ 1 }\] divide the numerator and denominator by 2 and solve

OpenStudy (anonymous):

Okay thanks!@Surjithayer

OpenStudy (anonymous):

\[\sqrt{3}\cot(\theta)-1=0\]\[\sqrt{3}\cot(\theta)=1\]\[\cot(\theta)=\frac{ 1 }{ \sqrt{3} }\]\[\tan(\theta)=\frac{ \sqrt{3} }{ 1 } \]\[\theta=\tan ^{-1}(\frac{ \sqrt{3} }{ 1 })\]\[\theta=\tan ^{-1}(\sqrt{3})\]

OpenStudy (anonymous):

Thanks so much😊@ Mateaus

OpenStudy (anonymous):

\[\tan \theta=\frac{ \frac{ \sqrt{3} }{ 2 } }{ \frac{ 1 }{ 2 } }=\tan \frac{ \pi }{ 3 }=\tan \left( n \pi+\frac{ \pi }{ 3 } \right)\] \[\theta=n \pi+\frac{ \pi }{ 3 }\] where n is an integer.

OpenStudy (thomas5267):

I think it is 60 deg and and 240 deg +- 360n deg.

OpenStudy (anonymous):

60 deg is correct answer